JosephGunnels:

f(x) = x2 + 1 g(x) = 5 – x (f + g)(x) = a. x2 + x – 4 b. x2 + x + 4 c. x2 – x + 6 d. x2 + x + 6

noahepaule:

@supie

supie:

Is this asking us to solve f(x) = x2 + 1 g(x) = 5 – x? I'm confused on the question.

MrMudd183:

Excuse me.... so what I would do is combine 1 & 5 to get 6, then put them in order it'll go like (f+g)(x)=x^2-x+6

JosephGunnels:

so what is the answer

Imagine:

This may help.

Imagine:

https://questioncove.com/study#/updates/5fc295d057639d2d6bb07c00 Finding the X Vaule.

JosephGunnels:

I am still confused

JosephGunnels:

thank u

MrMudd183:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @JosephGunnels I am still confused \(\color{#0cbb34}{\text{End of Quote}}\) All you need to do is combine the like terms and then merge the equations together

JosephGunnels:

ok